# Off forward non-SCHC contributions to exclusive vector quarkonium production from the “spin dependent BFKL Pomeron”

### v1

@@ I Introduction
The eikonal, forward BFKL Pomeron or unintegrated gluon distribution [1, 2, 3, 4] corresponds to the matrix element of a two-gluon $t$-channel exchange operator between incoming and outgoing light-cone proton states with the same helicity $\Lambda$. Indeed, in the eikonal (high-energy) and $t\to 0$ limits a helicity flip of the proton is not possible; the same applies to the transition of the photon to the vector meson state. This leads to so-called $s$-channel helicity conservation (SCHC) in eikonal exclusive $J/\psi$ or $\psi(2S)$ production when $t\to 0$ [5]. Specifically, in photoproduction or ultraperipheral proton-proton or nucleus-proton collisions, the vector meson (VM) is transversely polarized, and in fact its polarization $\bar{\lambda}$ is equal to the polarization $\lambda=\pm 1$ of the incoming photon.
In off-forward production, $-t>0$, non-SCHC corrections appear. The first kind preserves still the helicity of the proton, and is thus associated with the usual spin independent non-forward BFKL Pomeron [5, 6, 7] involving orbital angular momentum transfer to the $c-\bar{c}$ pair which scatters off the proton. Thus, the polarization of the $J/\psi$ may now be different from that of the photon, $\bar{\lambda}\neq\lambda$. In the limit of small $|t|\approx\bm{\Delta}_{\perp}^{2}$, non-SCHC contributions to the cross section in DIS (photoproduction) begin at linear (quadratic) order in the VM transverse momentum $\Delta_{\perp}$.
Here, we are concerned mainly with a second kind of non-SCHC corrections associated with a helicity flip of the proton. These are related to the so-called “spin-dependent Pomeron” (SDP) [8, 9, 10, 11]. Hard processes involve the partonic substructure of the proton, and orbital angular momentum may be transfered to the partons even in the eikonal limit, thereby allowing a helicity flip of the proton [^1] The SDP is associated with the $n=0$ BFKL eigenfunction with positive energy intercept.. The SDP is parameterized in terms of two distinct dipole-proton scattering amplitudes, associated with $(\bm{r}_{\perp}\cdot\bm{\Delta}_{\perp})(\bm{\epsilon}_{\perp}^{\Lambda}\cdot\bm{r}_{\perp})$ and $(\bm{\epsilon}_{\perp}^{\Lambda}\cdot\bm{\Delta}_{\perp})$ correlations, where $\bm{r}_{\perp}$ is the dipole vector and $\bm{\epsilon}_{\perp}^{\Lambda}=-(\Lambda,\mathrm{i})/\sqrt{2}$ the helicity vector, for proton helicity $\Lambda=\pm 1$. Therefore, the SDPs contribute at leading non-SCHC power ${\cal O}(\Delta_{\perp}^{2})$ to the cross section for exclusive vector quarkonium photoproduction (at small momentum transfer).
@@ III Spin-independent Pomeron exchange
$\begin{split}&\mathcal{M}^{\epsilon}_{\lambda\Lambda;-\lambda\Lambda}=4\pi eq_{c}N_{c}\int_{z}\int r_{\perp}\mathrm{d}r_{\perp}\mathcal{P}_{\epsilon}(r_{\perp},\Delta_{\perp})\mathcal{A}^{f}_{TT}(r_{\perp},z)\left[J_{4}(r_{\perp}\delta_{\perp})+J_{0}(r_{\perp}\delta_{\perp})\right]\sim\Delta_{\perp}^{2}\,,\end{split}$ (9)
@@
$\mathcal{M}^{\epsilon}_{\lambda\Lambda;0\Lambda}=-4\pi\mathrm{i}eq_{c}N_{c}\lambda\int_{z}{\rm sign}(z-\bar{z})\int r_{\perp}\mathrm{d}r_{\perp}\mathcal{P}_{\epsilon}(r_{\perp},\Delta_{\perp})\mathcal{A}_{TL}(r_{\perp},z)\,\left[J_{1}(r_{\perp}\delta_{\perp})\mathrm{e}^{-\mathrm{i}\lambda\phi_{\Delta}}-J_{3}(r_{\perp}\delta_{\perp})\mathrm{e}^{3\lambda\mathrm{i}\phi_{\Delta}}\right]\sim\Delta_{\perp}^{3}\,.$ (11)
@@
$\begin{split}\mathcal{M}_{0\Lambda;0\Lambda}&=4\pi eq_{c}N_{c}\int_{z}\int r_{\perp}\mathrm{d}r_{\perp}\mathcal{P}_{0}(r_{\perp},\Delta_{\perp})\mathcal{A}_{LL}(r_{\perp},z)J_{0}(r_{\perp}\delta_{\perp})\sim\Delta_{\perp}^{0}\,,\\ \mathcal{M}^{\epsilon}_{0\Lambda;0\Lambda}&=-8\pi eq_{c}N_{c}\int_{z}\int r_{\perp}\mathrm{d}r_{\perp}\mathcal{P}_{\epsilon}(r_{\perp},\Delta_{\perp})\mathcal{A}_{LL}(r_{\perp},z)J_{2}(r_{\perp}\delta_{\perp})\sim\Delta_{\perp}^{4}\,,\\ \end{split}$ (14)
@@
$\begin{split}\mathcal{M}_{0\Lambda;\bar{\lambda}\Lambda}&=-4\pi\mathrm{i}eq_{c}N_{c}\bar{\lambda}\mathrm{e}^{-\mathrm{i}\bar{\lambda}\phi_{\Delta}}\int_{z}\int r_{\perp}\mathrm{d}r_{\perp}\mathcal{P}_{0}(r_{\perp},\Delta_{\perp})\mathcal{A}_{LT}(r_{\perp},z)\,{\rm sign}(z-\bar{z})J_{1}(r_{\perp}\delta_{\perp})\sim\Delta_{\perp}^{1}\,,\\ \mathcal{M}^{\epsilon}_{0\Lambda;\bar{\lambda}\Lambda}&=4\pi\mathrm{i}eq_{c}N_{c}\mathrm{e}^{-\mathrm{i}\bar{\lambda}\phi_{\Delta}}\int_{z}\int r_{\perp}\mathrm{d}r_{\perp}\mathcal{P}_{\epsilon}(r_{\perp},\Delta_{\perp})\mathcal{A}_{LT}(r_{\perp},z)\left(J_{2-\bar{\lambda}}(r_{\perp}\delta_{\perp})-J_{2+\bar{\lambda}}(r_{\perp}\delta_{\perp})\right)\sim\Delta_{\perp}^{3}\,,\end{split}$ (15)
@@ IV Spin-dependent Pomeron exchange, proton helicity flip
$\begin{split}\mathcal{M}_{0\Lambda;0,-\Lambda}&=-2\pi eq_{c}N_{c}\Lambda\mathrm{e}^{\mathrm{i}\Lambda\phi_{\Delta}}\int_{z}\int r_{\perp}\mathrm{d}r_{\perp}\mathcal{P}_{S0}(r_{\perp},\Delta_{\perp})\mathcal{A}_{LL}(r_{\perp},z)\left[J_{2}(r_{\perp}\delta_{\perp})-J_{0}(r_{\perp}\delta_{\perp})\right]\sim\Delta_{\perp}^{1}\,,\\ \mathcal{M}_{0\Lambda;\bar{\lambda},-\Lambda}&=-2\pi\mathrm{i}N_{c}eq_{c}\mathrm{e}^{\mathrm{i}(-\bar{\lambda}+\Lambda)\phi_{\Delta}}\int_{z}\mathrm{sgn}(z-\bar{z})\int r_{\perp}\mathrm{d}r_{\perp}\,\mathcal{A}_{LT}(r_{\perp},z)\mathcal{P}_{S0}(r_{\perp},\Delta_{\perp})\\ &\times\left[J_{-\bar{\lambda}+\Lambda+1}(r_{\perp}\delta_{\perp})-J_{-\bar{\lambda}+\Lambda-1}(r_{\perp}\delta_{\perp})\right]\sim\Delta_{\perp}^{2}\,,\\ \end{split}$ (19)
@@
However, in the high energy limit employed here, the corresponding amplitudes $\mathcal{M}_{nf}$ and $\mathcal{M}_{f}$ are real, and so single spin asymmetries vanish. The real part of the interference can be captured through the double spin asymmetry $\Delta\sigma(\lambda_{e}=+1,\bm{S}_{\perp})-\Delta\sigma(\lambda_{e}=-1,\bm{S}_{\perp})$ where the incoming electron beam has longitudinal polarization $\lambda_{e}$. The presence of an additional spin supplies another factor of ‘$\mathrm{i}$’ and so:
@@ V Estimate of the spin dependent Pomeron at moderately small $x$
given by
$\begin{split}G_{2,\Lambda\Lambda^{\prime}}(\bm{q}_{1\perp},\bm{q}_{2\perp})&=\int[\mathrm{d}x_{i}]\int[\mathrm{d}^{2}k_{i}]\sum\limits_{\{\lambda_{i}\}}\sum_{j}\Big{[}\Phi^{*}_{\Lambda^{\prime}}(\lambda_{i},x_{i},\bm{k}_{i(j)\perp})\Psi^{*}(x_{i},\bm{k}_{i(j)\perp})\\ &-\frac{1}{2}\sum_{j^{\prime}\neq j}\Phi^{*}_{\Lambda^{\prime}}(\lambda_{i},x_{i},\bm{k}_{i(jj^{\prime})\perp})\Psi^{*}(x_{i},\bm{k}_{i(jj^{\prime})\perp})\Big{]}\Phi_{\Lambda}(\lambda_{i},x_{i},\bm{k}_{i\perp})\Psi(x_{i},\bm{k}_{i\perp})\,,\end{split}$ (25)
with $\bm{k}_{i(j)\perp}=\bm{k}_{i\perp}+(x_{i}-\delta_{ij})(\bm{q}_{1\perp}+\bm{q}_{2\perp})$, $\bm{k}_{i(jj^{\prime})\perp}=\bm{k}_{i\perp}+x_{i}(\bm{q}_{1\perp}+\bm{q}_{2\perp})-\bm{q}_{1\perp}\delta_{ij}-\bm{q}_{2\perp}\delta_{ij^{\prime}}$. The explicit form of the integration measures over quark LC momentum fractions $x_{i}$ and transverse momenta $\bm{k}_{i\perp}$ is given in Appendix B. For the spatial wave function $\Psi(x_{i},\bm{k}_{i\perp})$ we employ a simple model due to Schlumpf [38, 39].
@@ Acknowledgements.
We thank Y. Hatta, Yu. Kovchegov, and K. Kumerički for useful discussions, and R. Venugopalan for stimulating ideas that initiated this work. A. D. acknowledges support by the DOE Office of Nuclear Physics through Grant DE-SC0002307, and the hospitality of the EIC Theory Institute at Brookhaven National Laboratory from Sept. 2024 to Jan. 2025 where most of this work was performed. S. B. is supported by the Croatian Science Foundation (HRZZ) no. 5332 (UIP-2019-04).
@@ Appendix A Photon-vector meson overlaps
$\begin{split}&A_{\lambda=\pm 1,\bar{\lambda}=0}(\bm{l}_{\perp},\bm{l}_{1\perp},z)=-4M_{V}{z\bar{z}}(z-\bar{z})(\bm{\epsilon}_{\perp}^{\lambda}\cdot\bm{l}_{\perp})\,,\\ &A_{\lambda=\pm 1,\bar{\lambda}=\pm 1}(\bm{l}_{\perp},\bm{l}_{1\perp},z)=2\left[(\bm{\epsilon}_{\perp}^{\lambda}\cdot\bm{\epsilon}_{\perp}^{\bar{\lambda}*})(\bm{l}_{\perp}\cdot\bm{l}_{1\perp}+m^{2})+(z-\bar{z})^{2}(\bm{\epsilon}_{\perp}^{\lambda}\cdot\bm{l}_{\perp})(\bm{\epsilon}_{\perp}^{\bar{\lambda}*}\cdot\bm{l}_{1\perp})-(\bm{\epsilon}_{\perp}^{\lambda}\cdot\bm{l}_{1\perp})(\bm{\epsilon}_{\perp}^{\bar{\lambda}*}\cdot\bm{l}_{\perp})\right]\,,\end{split}$ (30)
@@
$\begin{split}&\mathcal{A}_{TT}^{nf}(r_{\perp},z)=\frac{1}{\pi}\frac{1}{z\bar{z}}\left[-(z^{2}+\bar{z}^{2})\varepsilon K_{1}(\varepsilon r_{\perp})\frac{\partial\phi_{T}}{\partial r_{\perp}}+m_{c}^{2}K_{0}(\varepsilon r_{\perp})\phi_{T}(r_{\perp},z)\right]\,,\\ &\mathcal{A}_{TT}^{f}(r_{\perp},z)=\frac{2}{\pi}\varepsilon K_{1}(\varepsilon r_{\perp})\frac{\partial\phi_{T}}{\partial r_{\perp}}\,,\end{split}$ (32)
@@
$\begin{split}&\mathcal{A}_{\lambda=0,\bar{\lambda}=0}(\bm{r}_{\perp},\bm{\Delta}_{\perp})=eq_{c}\int_{z}\mathrm{e}^{-\mathrm{i}\bm{\delta}_{\perp}\cdot\bm{r}_{\perp}}\mathcal{A}_{LL}(r_{\perp},z)\,,\\ &\mathcal{A}_{\lambda=0,\bar{\lambda}=\pm 1}(\bm{r}_{\perp},\bm{\Delta}_{\perp})=eq_{c}\bar{\lambda}\mathrm{e}^{-\mathrm{i}\bar{\lambda}\phi_{r}}\int_{z}\mathrm{e}^{-\mathrm{i}\bm{\delta}_{\perp}\cdot\bm{r}_{\perp}}\mathcal{A}_{LT}(r_{\perp},z)\,,\end{split}$ (35)
@@ Appendix B Pomeron amplitude at moderately small $x$ from a light-front constituent quark model
$[\mathrm{d}x_{i}]=\prod_{i=1\cdots 3}\frac{\mathrm{d}x_{i}}{2x_{i}}\,\delta\left(1-\sum_{i}x_{i}\right)\ ,\ \ \left[\mathrm{d}^{2}\bm{k}_{i\perp}\right]=\prod_{i=1\cdots 3}\frac{\mathrm{d}^{2}\bm{k}_{i\perp}}{(2\pi)^{3}}\,(2\pi)^{3}\,\delta^{(2)}\left(\sum_{i}\bm{k}_{i\perp}\right)~{}.$ (38)
@@
The quark mass $m_{q}=0.26$ GeV and the parameter $\beta=0.55$ GeV have been tuned to reproduce the electromagnetic “radius” and the anomalous magnetic moments of proton and neutron [^4] We have omitted the flavor wave function in (37) as this does not play a role in our analysis.; they also lead to reasonably good agreement of the Dirac and Pauli form factors of the proton with data. The normalization of the spatial wave function follows from the requirement that $K,\Lambda^{\prime}|P,\Lambda\right>=16\pi^{3}P^{+}\delta(P^{+}-K^{+})\delta^{(2)}(\bm{P}_{\perp}-\bm{K}_{\perp})\delta_{\Lambda\Lambda^{\prime}}$.
@@ Appendix C GTMDs of the eikonal dipole and the spin dependent Pomerons
$\begin{split}\int\mathrm{d}^{2}\bm{r}_{\perp}\mathrm{e}^{-\mathrm{i}\bm{k}_{\perp}\cdot\bm{r}_{\perp}}\mathcal{P}_{\Lambda\Lambda^{\prime}}(\bm{r}_{\perp},\bm{\Delta}_{\perp})&=-\frac{g^{2}(2\pi)^{3}}{4N_{c}M(\bm{k}_{\perp}^{2}-\bm{\Delta}_{\perp}^{2}/4)}\Bigg{[}M\delta_{\Lambda\Lambda^{\prime}}f_{1,1}(\bm{k}_{\perp},\bm{\Delta}_{\perp})+\Delta_{\perp}\frac{\bm{k}_{\perp}^{2}}{M^{2}}\delta_{\Lambda,-\Lambda^{\prime}}\Lambda\mathrm{e}^{\mathrm{i}\Lambda\phi_{k}}f_{1,2}(\bm{k}_{\perp},\bm{\Delta}_{\perp})\\ &+\Delta_{\perp}\Lambda\delta_{\Lambda,-\Lambda^{\prime}}\mathrm{e}^{\mathrm{i}\Lambda\phi_{\Delta}}\left(f_{1,3}(\bm{k}_{\perp},\bm{\Delta}_{\perp})-\frac{1}{2}f_{1,1}(\bm{k}_{\perp},\bm{\Delta}_{\perp})\right)\Bigg{]}\,,\end{split}$ (40)

### v2

@@ I Introduction
The eikonal, forward BFKL Pomeron or unintegrated gluon distribution [1, 2, 3, 4] corresponds to the matrix element of a two-gluon $t$-channel exchange operator between incoming and outgoing light-cone proton states with the same helicity $\lambda_{p}=\pm 1/2$. Indeed, in the eikonal (high-energy) and $t\to 0$ limits a helicity flip of the proton is not possible; the same applies to the transition of the photon to the vector meson state. This leads to so-called $s$-channel helicity conservation (SCHC) in eikonal exclusive $J/\psi$ or $\psi(2S)$ production when $t\to 0$ [5]. Specifically, in photoproduction or ultraperipheral proton-proton or nucleus-proton collisions, the vector meson (VM) is transversely polarized, and in fact its polarization $\bar{\lambda}$ is equal to the polarization $\lambda=\pm 1$ of the incoming photon.
In off-forward production, t>0-t>0, non-SCHC corrections appear. The first kind preserves still the helicity of the proton, and is thus associated with the usual spin independent non-forward BFKL Pomeron [5, 6, 7] involving orbital angular momentum transfer to the $c-\bar{c}$ pair which scatters off the proton. Thus, the polarization of the $J/\psi$ may now be different from that of the photon, $\bar{\lambda}\neq\lambda$. In the limit of small $|t|\approx\bm{\Delta}_{\perp}^{2}$, non-SCHC contributions to the cross section in DIS (photoproduction) begin at linear (quadratic) order in the VM transverse momentum $\Delta_{\perp}$.
Here, we are concerned mainly with a second kind of non-SCHC corrections associated with a helicity flip of the proton. These are related to the so-called “spin-dependent Pomeron” (SDP) [8, 9, 10, 11]. Hard processes involve the partonic substructure of the proton, and orbital angular momentum may be transfered to the partons even in the eikonal limit, thereby allowing a helicity flip of the proton [^1] The SDP is associated with the $n=0$ BFKL eigenfunction with positive energy intercept.. The SDP is parameterized in terms of two distinct dipole-proton scattering amplitudes, associated with $(\bm{r}_{\perp}\cdot\bm{\Delta}_{\perp})(\bm{\epsilon}_{\perp}^{\Lambda}\cdot\bm{r}_{\perp})$ and $(\bm{\epsilon}_{\perp}^{\Lambda}\cdot\bm{\Delta}_{\perp})$ correlations, where $\bm{r}_{\perp}$ is the dipole vector and $\bm{\epsilon}_{\perp}^{\Lambda}=-(\Lambda,\mathrm{i})/\sqrt{2}$ the helicity vector; hereafter we conventionally normalize the proton helicity defining $\Lambda={\color[rgb]{0,0,0}2\lambda_{p}}=\pm 1$. Therefore, the SDPs contribute at leading non-SCHC power ${\cal O}(\Delta_{\perp}^{2})$ to the cross section for exclusive vector quarkonium photoproduction (at small momentum transfer).
@@ III Spin-independent Pomeron exchange
[…]
@@
$\mathcal{M}^{\epsilon}_{\lambda\Lambda;0\Lambda}=-4\pi\mathrm{i}eq_{c}N_{c}\lambda{\color[rgb]{0,0,0}\mathrm{e}^{\mathrm{i}\lambda\phi_{\Delta}}}\int_{z}{\rm sign}(z-\bar{z})\int r_{\perp}\mathrm{d}r_{\perp}\mathcal{P}_{\epsilon}(r_{\perp},\Delta_{\perp})\mathcal{A}_{TL}(r_{\perp},z)\,\left[J_{1}(r_{\perp}\delta_{\perp})-J_{3}(r_{\perp}\delta_{\perp})\right]\sim\Delta_{\perp}^{3}\,.$ (11)
@@
[…]
@@
[…]
@@ IV Spin-dependent Pomeron exchange, proton helicity flip
[…]
@@
However, in the high energy limit employed here, the corresponding amplitudes $\mathcal{M}_{nf}$ and $\mathcal{M}_{f}$ are real, and so single spin asymmetries vanish. The real part of the interference can be captured through the double spin asymmetry $\Delta\sigma(\lambda_{e}=+1,\bm{S}_{\perp})-\Delta\sigma(\lambda_{e}=-1,\bm{S}_{\perp})$ where the incoming electron beam has longitudinal polarization $\lambda_{e}$. The presence of an additional spin supplies another factor of $\mathrm{i}$’ and so:
@@ V Estimate of the spin dependent Pomeron at moderately small $x$
$\langle K,\Lambda^{\prime}|\,J^{+a}(\bm{q}_{1\perp})J^{+b}(\bm{q}_{2\perp})\,|P,\Lambda\rangle=\frac{1}{2}\delta^{ab}\,\,16\pi^{3}\,P^{+}\,\delta(P^{+}-K^{+})\,\delta^{(2)}(\bm{P}_{\perp}-\bm{K}_{\perp}-\bm{q}_{1\perp}-\bm{q}_{2\perp})\,\,G_{2,\Lambda\Lambda^{\prime}}(\bm{q}_{1\perp},\bm{q}_{2\perp})\,$, (24)
given by
@@
with $\bm{k}_{i(j)\perp}=\bm{k}_{i\perp}+(x_{i}-\delta_{ij})(\bm{q}_{1\perp}+\bm{q}_{2\perp})$, $\bm{k}_{i(jj^{\prime})\perp}=\bm{k}_{i\perp}+x_{i}(\bm{q}_{1\perp}+\bm{q}_{2\perp})-\bm{q}_{1\perp}\delta_{ij}-\bm{q}_{2\perp}\delta_{ij^{\prime}}$. The explicit form of the integration measures over quark LC momentum fractions $x_{i}$ and transverse momenta $\bm{k}_{i\perp}$ is given in Appendix B. For the spatial wave function $\Psi(x_{i},\bm{k}_{i\perp})$ we employ a simple model due to Schlumpf [38, 39]. The spin flip two-gluon exchange amplitude has already been studied long ago in a non-relativistic quark-diquark model of the proton [54, 55]. It was found that in the soft regime the spin–flip coupling is small and that it is quite sensitive to diquark correlations. Here, we restrict to hard gluon exchanges where $|\bm{q}_{1\perp}|,|\bm{q}_{2\perp}|\gg\Lambda_{\mathrm{QCD}}$ due to the fact that the $t$-channel gluons also couple to a small dipole. In this regime diquark correlations in the proton may be less significant although such questions would deserve further scrutiny in the future.
@@ Acknowledgements.
We thank Y. Hatta, Yu. Kovchegov, and K. Kumerički for useful discussions, and R. Venugopalan for stimulating ideas that initiated this work. We thank M. Spinelli for pointing out an error in an earlier version of this paper. A. D. acknowledges support by the DOE Office of Nuclear Physics through Grant DE-SC0002307, and the hospitality of the EIC Theory Institute at Brookhaven National Laboratory from Sept. 2024 to Jan. 2025 where most of this work was performed. S. B. is supported by the Croatian Science Foundation (HRZZ) no. 5332 (UIP-2019-04).
@@ Appendix A Photon-vector meson overlaps
[…]
@@
[…]
@@
[…]
@@ Appendix B Pomeron amplitude at moderately small $x$ from a light-front constituent quark model
$\mathrm{d}x_{i}]=\prod_{i=1\cdots 3}\frac{\mathrm{d}x_{i}}{2x_{i}}\,\delta\left(1-\sum_{i}x_{i}\right)\ ,\ \ \left[\mathrm{d}^{2}\bm{k}_{i\perp}\right]=\prod_{i=1\cdots 3}\frac{\mathrm{d}^{2}\bm{k}_{i\perp}}{(2\pi)^{3}}\,(2\pi)^{3}\,\delta^{(2)}\left(\sum_{i}\bm{k}_{i\perp}\right)~.$ (38)
@@
The quark mass $m_{q}=0.26$ GeV and the parameter $\beta=0.55$ GeV have been tuned to reproduce the electromagnetic “radius” and the anomalous magnetic moments of proton and neutron [^4] We have omitted the flavor wave function in (37) as this does not play a role in our analysis.; they also lead to reasonably good agreement of the Dirac and Pauli form factors of the proton with data. The normalization of the spatial wave function follows from the requirement that $\left<K,\Lambda^{\prime}|P,\Lambda\right>=16\pi^{3}P^{+}\delta(P^{+}-K^{+})\delta^{(2)}(\bm{P}_{\perp}-\bm{K}_{\perp})\delta_{\Lambda\Lambda^{\prime}}$.
@@ Appendix C GTMDs of the eikonal dipole and the spin dependent Pomerons
[…]